4ab5 is a four digit number divisible by 55 where a, b are unknown digits. Then b – a is :

Grade:6

2 Answers

Sourabh Singh IIT Patna
7 years ago
hii
It is very easy as he number given is a multiple of 11*5 so it must follow he divsibility rule of 11 and 5

use that and you will get the relation in a and b which will give you answer
Yuvraj Singh
5 years ago
55=5*11.
We know in the no. 4ab5 last digit is 5 so its divisible by 5.
Its divisible by 11 also if its divisible by 11 then the sum of its odd and even terms must be divisible by
11. 
4+b-(a+5)=a no. Divisible by 11.
b-a=(a no. Divisible by 11 ) +1.
now since b and a are single digit no. so 0 is the only no. which satisfies the equation. So the answer is 1.